# Bounds to the Basset-Boussinesq force on particle laden stratified flows

### v1

@@ Abstract
The Basset-Boussinesq force is often perfunctory neglected when studying small inertial particles in turbulence. The force results from the diffusion of vorticity from the particles, and as it depends on the particles past history, it complicates the dynamics by turning their equations of motion into integro-differential equations. However, this force is of the same order as other viscous forces acting on the particles, and beyond convenience, the reasons for neglecting it are unclear. We derive strict bounds for the magnitude of the Basset-Boussinesq force in stably stratified flows, in contexts of interest for geophysical turbulence. The bounds are validated by direct numerical simulations. The Basset-Boussinesq force can be neglected when a buoyancy Stokes number $\textrm{Sb}=N\tau_{p}$ is small, where $N$ is the flow Brunt-Väisälä frequency and $\tau_{p}$ is the particles Stokes time. For sufficiently strong stratification, or particles with large inertia, this force must be considered in the dynamics.
@@ I Introduction
The dynamics of inertial particles submerged in turbulent flows plays a crucial role in various geophysical contexts, ranging from phenomena in coastal environments and lakes to the atmosphere and the oceans [1, 2, 3, 4, 5]. Despite extensive research on the influence of turbulence on the transport and spatial distribution of particles, a complete description of their dynamics in geophysical systems remains elusive. Furthermore, the complexity extends to the environmental sphere, where the dynamics of plankton and algae, and the presence of microplastics in the oceans, introduce additional layers of difficulty to particle dispersion, and emphasize the urgent need to understand these phenomena [6]. In many of these cases the complexity of modeling and simulating these systems imposes the need to use reduced models that simplify the physics. Recently, both experiments [7, 8] and particle-resolved simulations [9, 10] have provided valuable insights into particle dynamics in different regimes. Many hydrodynamic forces operate over the particles at small scales within turbulent flows, influencing their interactions, collisions, and cluster formation [9].
@@
The modeling of geophysical flows, particularly those featuring stable density stratification, presents also distinct challenges even in the absence of particles. Anisotropy characterizes stably stratified turbulence, setting it apart from homogeneous isotropic turbulence (HIT) [11, 12, 13, 14, 15]. Within these flows, the impact of stratification is evident in the reduction of vertical velocity, leading to confined, quasi-horizontal layered motions, and to the generation of vertically sheared horizontal winds (VSHWs) marked by significant vertical variability [16]. The stratified environment introduces a restoring force, allowing for the coexistence of waves with turbulence, each exhibiting a distinct spectral scaling compared to HIT. Stably stratified turbulence manifests an anisotropic inertial subrange, fostering a direct energy cascade between buoyancy and Ozmidov scales [17, 18]. Notably, studies also suggest that larger-scale quasi-horizontal motions serve as a continuous source of small-scale turbulence, provided the local Reynolds number does not fall below a critical threshold [11]. Examining these phenomena, it becomes apparent that in stably stratified turbulence, as described by Herring and Métais [19] and by Riley and Lelong [20], thin layers of large quasihorizontal vortical structures coexist with internal gravity waves [21]. All these features underscore the complex interplay of forces and structures characterizing turbulence in geophysical flows.
@@
In this context, several recent studies have considered the dynamics of different particles advected by stable stratified flows. As a first example, analysis of Lagrangian tracers in stratified turbulence was performed in [22, 23]. When particles have inertia, for small-sized particles in a turbulent flow the Maxey-Riley approximation provides a set of equations to describe their dynamics [24]. This approximation was extended to stratified flows for heavy [25] and neutrally buoyant and light particles [25, 26, 27]. As particles increase in size, Faxen corrections proportional to the square radius of the particles become relevant [25]. But even for very small particles the Maxey-Riley equation is an integro-differential system for the particles evolution that depends on the particles past history. To further simplify the problem the Basset-Boussinesq force, resulting from the diffusion of vorticity away from the particles along their trajectories, is often neglected. Methods to reduce the computational cost to estimate this force were devised [28], and a few studies have considered its effect in particle laden stratified flows [25]. It was seen that this force becomes particularly relevant in the presence of flow stratification, as particles experience significant acceleration and deceleration resulting from oscillatory motions caused by bouyancy. However, it is still unclear under what general conditions this force becomes dominant, or negligible.
@@
In this work we derive a bound to the Basset-Bousinesq force in a stratified fluid, that allows the estimation of under which conditions this force can be neglected. To estimate this bound we consider the Maxey-Riley equation for small inertial particles in a stably stratified fluid. We also conduct direct numerical simulations of the Boussinesq equations for the fluid, along with the Maxey-Riley equation with and without the Basset-Bousinesq (or history) term. We compare particle dispersion and particle velocity statistics, as well as the particle preferential clustering, for different values of the controlling parameters. We show that the bound allows estimation of under what conditions neglecting the Basset-Boussinesq history term does not affecting the statistical behavior of any of these observables.
@@ II Equations of motion
We consider a stably stratified fluid described by the incompressible Boussinesq equations for the velocity ${\bf u}=(u,v,w)$ and for mass density fluctuations $\rho^{\prime}$,
@@
where $p$ is the correction to the hydrostatic pressure, $\nu$ is the kinematic viscosity, ${\bf f}$ is an external mechanical forcing, $N$ is the Brunt-Väisälä frequency (which in this approximation sets the stratification), and $\kappa$ is the diffusivity. In terms of the background density gradient, the Brunt-Väisälä frequency is $N^{2}=-(g/\rho_{0})(d\bar{\rho}/dz)$, with $d\bar{\rho}/dz$ the imposed linear background stratification, and $\rho_{0}$ the mean fluid density. We write scaled density fluctuations $\zeta$ in units of velocity by defining $\zeta=g\rho^{\prime}/(\rho_{0}N)$. All quantities are then made dimensionless using a characteristic length $L_{0}$ and a characteristic velocity $U_{0}$ in the domain, resulting in
@@
We describe small inertial particles using the Maxey-Riley equation, under an approximation consistent with those made to derive the Boussinesq equations [27]. Under the Boussinesq approximation for a stratified fluid, Eqs. (4) and (5) are obtained from the Navier-Stokes equations after neglecting density fluctuations everywhere except in the buoyancy force. Thus, for the dynamics of the particles we also approximate the density and the mass of the fluid displaced by the particles by their mean values, i.e., $\rho_{f}\approx\bar{\rho}_{f}=\rho_{0}$ and $m_{f}\approx\bar{m}_{f}=\rho_{0}V_{p}$ (where $V_{p}$ is particle volume), except in the gravity term. In that term we keep the full fluid density dependence, $\rho_{f}=\rho_{0}+(d\bar{\rho}/dz)(z-z_{0})+\rho^{\prime}$ (note that for a stably stratified fluid $d\bar{\rho}/dz<0$). For simplicity we also assume that the typical length over which the velocity field changes appreciably is much larger than the particle radius $a$, and Faxén corrections are thus neglected. With these approximations the equation for the particles is
@@
where ${\bf x}_{p}(t)$ is the particle position, ${\bf v}$ is the particle velocity, ${\bf u}({\bf x}_{p},t)$ is the fluid velocity at the particle position, $D/Dt$ is the Lagrangian derivative, $\textrm{d}/\textrm{d}t$ is the time derivative following the particle trajectory, $z_{0}$ is the height at which particles are neutrally buoyant, and $\tau_{p}=(m_{p}+\bar{m}_{f}/2)/(6\pi a\bar{\rho}_{f}\nu)$ is the particle relaxation time. For a spherical particle $\tau_{p}=a^{2}/(3\nu)$ when $\bar{m}_{f}/m_{p}=1$ (under these approximations, any other choice for $\bar{m}_{f}/m_{p}$ is equivalent to changing the reference value $\rho_{0}$ and results in particles being neutrally buoyant at a different height, which can be absorbed into a displacement of $z_{0}$). We can finally define the particles Stokes number as $\textrm{St}=\tau_{p}/\tau_{\eta}$, where $\tau_{\eta}=(\nu/\varepsilon)^{1/2}$ is the Kolmogorov time scale, and $\varepsilon$ is the fluid kinetic energy dissipation rate.
@@ III Bounds to the Basset-Boussinesq force
Note that when the Faxén corrections are neglected, $\textbf{h}(t)$ is the slip velocity of each particle. Both the Stokes term in Eq. (6), $\textbf{h}(t)/\tau_{p}$, and the Basset-Boussinesq history term, depend on the particle inertia. Thus, it is reasonable to bound the Basset-Boussinesq force compared against the Stokes force.
In the most general (not stratified) case, Van Hinsberg et al. [28] proved that the Basset-Boussinesq force is finite at any given time, provided some constraints on $\textbf{f}(t)$ and $\textbf{g}(t)$ are satisfied. First, $\textbf{f}(t)$ must be a continuous function, and its derivative must exist almost everywhere. Second, the infinity norm of $\textbf{f}(t)$ and $\textbf{g}(t)$ must be finite. These constraints on $\textbf{f}(t)$ and $\textbf{g}(t)$ are
$\textbf{f}\in C^{0},~{}~{}~{}~{}~{}\|\textbf{f}\|_{\infty}=B_{1},~{}~{}~{}~{}~{}\|\textbf{g}\|_{\infty}=B_{2},$ (8)
@@
$\|\textbf{f}\|_{\infty}=\inf\{C\geq 0:|\textbf{f}(t)|\leq C\mbox{ almost anywhere}\},$ (9)
where ||$\cdot$|| ⋅ | represents the typical vector length. Let’s assume that these conditions are fulfilled for particles in turbulent flows with $\textbf{f}(t)={\bf v}_{\textrm{slip}}(t)$. Under these conditions an upper bound for the Basset-Boussinesq force $\textbf{F}_{\textrm{BB}}$ can be obtained. The convolution is divided into two parts to control both the singularity in the kernel and in the tail of the integral [28]:
@@
where $c_{\textrm{BB}}=\sqrt{3/(\pi\tau_{p})}$, and where this bound holds for all times.
In the stratified case, note that Eq. (6) can be rewritten as
@@
where ${\bf F}(t)$ is a forcing acting on each particle. In the vertical component this is an equation for a driven damped oscillator with frequency $\sqrt{2/3}N$ and with damping constant $1/(2\tau_{p})$ [27]. As in the vertical direction we have fast fluctuations of the particles velocities, it makes sense to continue looking at this component of the equation. In [27] it was shown that assuming that in a stratified flow the vertical fluid velocity is dominated by internal gravity waves, and thus approximating fluid element displacements as those resulting from the propagation of these waves, $z_{f}=z_{0}+\zeta/N$, $\dot{z}_{f}=w({\bf x}_{p}(t),t)$, and $\ddot{z}_{f}=Dw/Dt$, we can further write the vertical component of Eq. (6) as
@@
The homogeneous solution to this equation, neglecting damping (as we are looking for maximum bounds for the velocities and viscous forces acting on the particle), is $z_{p}-z_{f}=A_{0}\exp(i\sqrt{2/3}Nt)$, and thus we can approximate the vertical component of the slip velocity as $w_{\textrm{slip}}=W_{0}\exp(i\sqrt{2/3}Nt)$ (real parts are assumed everywhere). The vertical component of the Stokes force can then be written as,
@@
where we used that the vertical component of convolution function in the history term is $g_{z}(t)=\textrm{d}w_{\textrm{slip}}/\textrm{d}t=i\sqrt{2/3}Nw_{\textrm{slip}}(t)$.
@@
which in the following we will call the buoyancy Stokes number, as it corresponds to a ratio between the particle response time and the Brunt-Väisälä period. We expect that for sufficiently small Sb the effect of the Basset-Boussinesq force should be negligible.

### v2

@@ Abstract
The Basset-Boussinesq force is often perfunctorily neglected when studying small inertial particles in turbulence. This force arises from the diffusion of vorticity from the particles and, since it depends on the particles’ past history, complicates the dynamics by transforming their equations of motion into integro-differential equations. However, this force is of the same order as other viscous forces acting on the particles, and beyond convenience, the reasons for neglecting it are unclear. This study addresses the following question: Under what conditions can the Basset-Boussinesq force be neglected in light particles in geophysical flows? We derive strict bounds for the magnitude of the Basset-Boussinesq force in stably stratified flows, in contexts of interest for geophysical turbulence. The bounds are validated by direct numerical simulations. The Basset-Boussinesq force is negligible when a buoyancy Stokes number $\textrm{Sb}=N\tau_{p}$ is small, where $N$ is the flow Brunt-Väisälä frequency and $\tau_{p}$ is the particle’s Stokes time. Interestingly, for most oceanic particles this force may be negligible. Only for very strong stratification, or for particles with very large inertia, this force must be considered in the dynamics.
@@ I Introduction
The dynamics of inertial particles submerged in turbulent flows plays a crucial role in various geophysical contexts, from phenomena in coastal environments and lakes, to the atmosphere and the oceans [1, 2, 3, 4, 5]. Despite extensive research on the influence of turbulence on the transport and spatial distribution of particles, a complete description of their dynamics in geophysical systems remains elusive. Furthermore, the complexity increases in the environmental sphere, where the dynamics of plankton and algae, and the presence of microplastics in the oceans, introduce additional layers of difficulty to particle dispersion, and emphasize the urgent need to understand these phenomena [6]. In many of these cases the complexity of modeling and simulating these systems imposes the need to use reduced models that simplify the physics. Recently, both experiments [7, 8] and particle-resolved simulations [9, 10] have provided valuable insights into particle dynamics in different regimes. Many hydrodynamic forces operate over the particles at small scales within turbulent flows, influencing their interactions, collisions, and cluster formation [9].
@@
The modeling of geophysical flows, particularly those featuring stable density stratification, presents also distinct challenges even in the absence of particles. Anisotropy characterizes stably stratified turbulence, setting it apart from homogeneous isotropic turbulence (HIT) [11, 12, 13, 14, 15]. Within these flows, the impact of stratification is evident in the reduction of vertical velocity, leading to confined, quasi-horizontal layered motions, and to the generation of vertically sheared horizontal winds marked by significant vertical variability [16]. The stratified environment introduces a restoring force, allowing for the coexistence of waves with turbulence, each exhibiting a distinct spectral scaling compared to HIT. Stably stratified turbulence manifests an anisotropic inertial subrange, fostering a direct energy cascade between buoyancy and Ozmidov scales [17, 18]. Notably, studies also suggest that larger-scale quasi-horizontal motions serve as a continuous source of small-scale turbulence, provided the local Reynolds number does not fall below a critical threshold [11]. Examining these phenomena, it becomes apparent that in stably stratified turbulence, as described by Herring and Métais [19] and by Riley and Lelong [20], thin layers of large quasihorizontal vortical structures coexist with internal gravity waves [21]. All these features underscore the complex interplay of forces and structures characterizing turbulence in geophysical flows.
@@
In this context, several recent studies have considered the dynamics of different particles advected by stable stratified flows. As a first example, analysis of Lagrangian tracers in stratified turbulence was performed in [22, 23]. When particles have inertia, for small-sized particles in a turbulent flow the Maxey-Riley-Gatignol approximation provides a set of equations to describe their dynamics [24, 25]. This approximation was either used heuristically, or extended to stratified flows, for heavy [26] and for neutrally buoyant and light particles [26, 27, 28] (see [28] for a derivation of equations for inertial particles compatible with the Boussinesq approximation in stratified fluids). As particles increase in size, Faxén corrections proportional to the square radius of the particles become relevant [26]. But even for very small particles the Maxey-Riley-Gatignol equation is an integro-differential system for the particles’ evolution that depends on the particles’ past history. To further simplify the problem the Basset-Boussinesq force, resulting from the diffusion of vorticity away from the particles along their trajectories, is often neglected. Methods to reduce the computational cost to estimate this force were devised [29], and a few studies have considered its effect in particle laden stratified flows [26]. It was seen that this force becomes particularly relevant in the presence of flow stratification, as particles experience significant acceleration and deceleration resulting from oscillatory motions caused by buoyancy. However, it is still unclear under what general conditions this force becomes dominant, or negligible.
As an example of an application of the study of particle-laden flows in geophysical contexts, we can mention models based on heuristically modified Maxey-Riley-Gatignol equations, such as the one used by Beron-Vera et al. for the transport of Sargassum, marine debris, and garbage patches in the ocean [30, 31, 32]. The authors modify the Maxey-Riley-Gatignol equation to account for the effect of the Coriolis force and oceanic vorticity, leading to a specific formulation for the evolution of inertial particles that incorporates these effects. Another example is provided by the work in [33], which empirically models the dynamics of floaters, such as plankton in the ocean, by considering fluid stratification while neglecting rotation. Both models highlight the importance of considering inertial effects and drag forces, as well as environmental forces, in the accurate description of particle dynamics in geophysical flows [34]. But both models also neglect the Basset-Boussinesq history force.
@@
In this work we derive a bound to the Basset-Boussinesq force in a stratified fluid, that allows estimation of conditions under which this force can be neglected. To obtain this bound we consider the Maxey-Riley-Gatignol equation for small inertial particles in a stably stratified fluid (see Sec. II, where we also discuss its range of applicability and typical values of parameters in geophysical flows, and Sec. III where we present the bound). We also conduct direct numerical simulations of the Navier-Stokes equation under the Boussinesq approximation for the fluid, along with the Maxey-Riley-Gatignol equation with and without the Basset-Boussinesq (or history) term (see Sec. IV). In Sec. V we compare particle dispersion and particle velocity statistics, as well as the particle preferential clustering, for different values of the controlling parameters. We show that the bound allows estimation of under what conditions neglecting the Basset-Boussinesq history term does not affect the statistical behavior of any of these observables (see Sec. VI for the main conclusions).
@@ II Equations of motion
We consider a stably stratified fluid described by the incompressible Navier-Stokes equation under the Boussinesq approximation for the velocity ${\bf u}=(u,v,w)$, and an equation for mass density fluctuations $\rho^{\prime}$,
@@
where $p$ is the correction to the hydrostatic pressure, $\nu$ is the kinematic viscosity, ${\bf f}$ is an external mechanical forcing, $N$ is the aforementioned Brunt-Väisälä frequency (which in this approximation sets the stratification), and $\kappa$ is the diffusivity. In terms of the background density gradient, the Brunt-Väisälä frequency is $N^{2}=-(g/\rho_{0})(d\bar{\rho}/dz)$, with $d\bar{\rho}/dz$ the imposed linear background stratification, and $\rho_{0}$ the mean fluid density. We write scaled density fluctuations $\zeta$ in units of velocity by defining $\zeta=g\rho^{\prime}/(\rho_{0}N)$. All quantities are then made dimensionless using a characteristic length $L_{0}$ and a characteristic velocity $U_{0}$ in the domain, resulting in
@@
We describe small neutrally buoyant inertial particles using the Maxey-Riley-Gatignol equation, under an approximation consistent with those made in the Boussinesq approximation used to describe the fluid (see [28] for a derivation assuming that particles are sufficiently small and thus that the flow in the vicinity of the particles is homogeneous). Note that under the Boussinesq approximation for a stratified fluid, Eqs. (4) and (5) are obtained from the Navier-Stokes equations after neglecting density fluctuations everywhere except in the buoyancy (or gravity) force. Thus, for the dynamics of the particles we also approximate the density and the mass of the fluid displaced by the particles by their mean values, i.e., $\rho_{f}\approx\bar{\rho}_{f}=\rho_{0}$ and $m_{f}\approx\bar{m}_{f}=\rho_{0}V_{p}$ (where $V_{p}$ is particle volume), except in the gravity term. In that term we keep the complete fluid density dependence, $\rho_{f}=\rho_{0}+(d\bar{\rho}/dz)(z-z_{0})+\rho^{\prime}$ (note that for a stably stratified fluid $d\bar{\rho}/dz<0$). Therefore, the full effect of gravity and buoyancy forces acting on the particles is preserved. For simplicity we also assume that the typical length over which the velocity field changes appreciably is much larger than the particle radius $a$, and Faxén corrections are thus neglected. With these approximations the equation for the neutrally buoyant particles is
@@
where ${\bf x}_{p}(t)$ is the particle position, ${\bf v}(t)$ is the particle velocity, ${\bf u}({\bf x}_{p},t)$ is the fluid velocity at the particle position, $D/Dt$ is the Lagrangian derivative, $\textrm{d}/\textrm{d}t$ is the time derivative following the particle trajectory, $z_{0}$ is the height at which particles are neutrally buoyant, $\zeta$ is the fluid density fluctuation at the particle position, and $\tau_{p}=(m_{p}+\bar{m}_{f}/2)/(6\pi a\bar{\rho}_{f}\nu)$ is the particle relaxation time. For a spherical particle $\tau_{p}=a^{2}/(3\nu)$ when $\bar{m}_{f}/m_{p}=1$ (note that this choice does not make the dynamics independent of the ratio of densities of the fluid and the particles; under these approximations any value of $\bar{m}_{f}/m_{p}\neq 1$ is equivalent to changing the reference value $\rho_{0}$ of the fluid, and results in particles being neutrally buoyant at a different height, which can in turn be absorbed into a new reference value of $z_{0}$).
The second term on the r.h.s. of Eq. (6) corresponds to gravity and buoyancy forces. Indeed, when multiplied by $m_{p}$, $-(2/3)\,m_{p}N^{2}\,(z_{p}-z_{0})=(2/3)\,m_{p}g\,(\rho_{0}^{-1}d\bar{\rho}/dz)(z_{p}-z_{0})$, and $(2/3)\,m_{p}N\zeta=(2/3)\,m_{p}g\,\rho^{\prime}/\rho_{0}$. Vertical forces acting on the particle thus depend on the height at which the particle is (the fluid background stratification, which is negative, pushes heavier particles downwards and lighter particles upwards), as well as on local density fluctuations of the fluid density (variations in the total fluid density associated with $\rho^{\prime}$ affect the buoyancy forces felt by the particles). Gravity and buoyancy can thus be the most important forces in the vertical dynamics of the particles. Note a similar role is played in the horizontal particles’ velocities in the heuristic model in [35, 36] by the Coriolis force.
We finally define the particles Stokes number as $\textrm{St}=\tau_{p}/\tau_{\eta}$, where $\tau_{\eta}=(\nu/\varepsilon)^{1/2}$ is the Kolmogorov time scale, and $\varepsilon$ is the fluid kinetic energy dissipation rate. With this definition we can discuss typical values of this number and of dimensional parameters of particles in geophysical contexts, which will be useful in Secs. III and IV. A typical value of the Brunt-Väisälä frequency in oceanic flows is $N\approx 10^{-3}$ s-1. The oceanic lower mixed layer or the upper seasonal thermocline have a mean Kolmogorov time scale $\tau_{\eta}\approx 3$ s, and Kolmogorov dissipation scales between $0.1$ and 10 mm [37]. Nanoplankton and microplankton have typical sizes between $2$ to $200$ microns [38], with Stokes numbers varying between $\textrm{St}\approx 10^{-5}$ to $10^{-1}$. Other neutrally buoyant particles in the ocean have particle radii between $1$ to $50$ mm, and particle response times between $\tau_{p}\approx 0.1$ to $100$ s [39], resulting in Stokes numbers between $\textrm{St}\approx 0.3$ to $30$. The mass ratio for oceanic neutrally buoyant particles can be as close to unity as $\approx 1.003$ [40]. Sargassum, seaweed, and marine debris in upper oceanic layers can have larger sizes, but are still modeled with modified Maxey-Riley-Gatignol equations using effective Stokes numbers smaller than 1 [35, 36]. Even though it is not the focus of this study, we also briefly mention that in the atmosphere the Brunt-Väisälä frequency varies between $0.1$ to $0.01$ s-1 depending on altitude and atmospheric conditions [41], while light particles can have response times on the order of $0.05$ s [42].
@@ III Bounds to the Basset-Boussinesq force
Note that $\textbf{h}(t)$ is the slip velocity of each particle only as long as Faxén corrections are neglected. Both the Stokes term in Eq. (6), $\textbf{h}(t)/\tau_{p}$, and the Basset-Boussinesq history term, depend on the particle inertia. Thus, it is reasonable to bound the Basset-Boussinesq force compared against the Stokes force.
In the most general (not stratified) case, Van Hinsberg et al. [29] proved that the Basset-Boussinesq force is finite at any given time, provided some constraints on $\textbf{h}(t)$ and $\textbf{g}(t)$ are satisfied. First, $\textbf{h}(t)$ must be a continuous function, and its derivative must exist almost everywhere. Second, the infinity norm of $\textbf{h}(t)$ and $\textbf{g}(t)$ must be finite. These constraints on $\textbf{h}(t)$ and $\textbf{g}(t)$ are
$\textbf{h}\in C^{0},~{}~{}~{}~{}~{}\|\textbf{h}\|_{\infty}=B_{1},~{}~{}~{}~{}~{}\|\textbf{g}\|_{\infty}=B_{2},$ (8)
@@
$\|\textbf{h}\|_{\infty}=\inf\{C\geq 0:|\textbf{h}(t)|\leq C\mbox{ almost anywhere}\},$ (9)
where ||$\cdot$|| ⋅ | represents the typical vector length. Let’s assume that these conditions are fulfilled for particles in turbulent flows with $\textbf{h}(t)={\bf v}_{\textrm{slip}}(t)$. Under these conditions an upper bound for the Basset-Boussinesq force $\textbf{F}_{\textrm{BB}}$ can be obtained. The convolution is separated into two parts to control both the singularity in the kernel and in the tail of the integral [29]:
@@
where $c_{\textrm{BB}}=\sqrt{3/(\pi\tau_{p})}$, and where this bound holds for all times (note that while this separation can also be used to implement a numerical scheme for the Basset-Boussinesq force, as in [43], this result uses the separation only to bound the force and does not depend on the method used to estimate the force numerically).
In the stratified case, Eq. (6) can be rewritten as
@@
where
[…]
is a force acting on each particle. This force ${\bf F}$ should not be confused with the external force ${\bf f}$ acting on the fluid in Eq. (4); note that Eq. (12) includes all the forces exerted on the particle by the fluid in Eq. (6).
In the vertical component, Eq. (11) is an equation for a driven damped oscillator with frequency $\sqrt{2/3}N$ and with damping constant $1/(2\tau_{p})$ [28]. As in the vertical direction we have fast fluctuations of particles velocities, it makes sense to continue looking at this component of the equation. In [28] it was shown that, assuming that in a stratified flow the vertical fluid velocity is dominated by internal gravity waves, and thus approximating fluid element displacements by those resulting from the propagation of these waves, $z_{f}=z_{0}+\zeta/N$, $\dot{z}_{f}=w({\bf x}_{p}(t),t)$, and $\ddot{z}_{f}=Dw/Dt$, we can write the vertical component of Eq. (6) as
@@
The vertical component of the first three terms in the force ${\bf F}$ are now absorbed on the l.h.s. of this equation, leaving only ${\bf F}_{\textrm{BB}}$ on the right. The homogeneous solution to this equation, neglecting damping (as we are looking for maximum bounds for the velocities and viscous forces acting on the particle), is $z_{p}-z_{f}=A_{0}\exp(i\sqrt{2/3}Nt)$, and thus we can approximate the vertical slip velocity as $w_{\textrm{slip}}=W_{0}\exp(i\sqrt{2/3}Nt)$ (real parts are assumed everywhere). The vertical component of the Stokes force can then be written as,
@@
where we used that the vertical component of the convolution function in the history term is $g_{z}(t)=\textrm{d}w_{\textrm{slip}}/\textrm{d}t=i\sqrt{2/3}Nw_{\textrm{slip}}(t)$.
@@
which in the following we will call the buoyancy Stokes number, as it corresponds to a ratio between the particle response time and the Brunt-Väisälä period  [^1] It is possible to show that $\textrm{Sb}\sim\textrm{St}/\textrm{Rb}^{\frac{1}{2}}$, which provides a measure of the particles’ inertia relative to the intensity of turbulence at the buoyancy scale.. We expect that for sufficiently small Sb the effect of the Basset-Boussinesq force should be negligible.
It is important to remark that the bound given by Eq. (19) can only be valid when the stratification of the fluid is sufficiently strong, and cannot hold in the limit of an isotropic and homogeneous fluid with $N\to 0$. This is a direct consequence of Eq. (11) being that of a driven damped oscillator only when the fluid is stratified. Later, when deriving Eq. (16), we also assumed vertical fluid velocity fluctuations are faster than horizontal velocity fluctuations, which amounts to assuming that the fluid Froude number $\textrm{Fr}=1/(TN)<1$, where $T$ is some characteristic fluid turnover time (see Sec. IV for detailed definitions).